Ә) Да Д - ТДse ГОСТ Кү : 24 12 ОН - (2 ч. 3s ) OS ора има . So - pH - ( 3 ota.s) as 24.7 =(( 31-%) 365) - Эч 35) Os Дs = Ана gibbs free equation please help me find delta s and delta h with the given delta gs
i set up my equation just unsure how to solve
delta h at 294.35 k: 24.7
delta h at 366.5 k : 32.86
Answer #1

Sol :-

Given equations are :  

24.7 = ΔH - (294.35) ΔS ................(1)

32.86 = ΔH - (366.5) ΔS ................(2)

On Subtracting both the equations, We have

24.7 - 32.86 = - 294.35 ΔS + 366.5 ΔS

- 8.16 = 72.15 ΔS

ΔS = -8.16 / 72.15

ΔS = - 0.113

Substitute the value of ΔS in equation (1)

24.7 = ΔH - (294.35) (- 0.113 )

24.7 = ΔH + 32.26

ΔH = 24.7 - 32.26

ΔH = - 7.56